# WaterMax: breaking the LLM watermark detectability-robustness-quality trade-off

### v1

@@ 4.1 Exploring the text space
4.2 Cumulative scores
$p=\mathbb{P}(S>s)=1-F_{S}(s;L).$ (3)
For the sake of clarity, suppose that the chunks are composed of $\ell$ tokens. At iteration $i$, the $j$-th candidate has a score denoted $s_{i,j}$ and $n$ drafts for its following chunk are proposed. This produces $n$ incremented scores per candidate, thus $mn$ variables: $s_{i,j}+\delta s_{i,j,k}$, $1\leq j\leq m$, $1\leq k\leq n$. Since these scores are converted into $p$-values by the same function, i.e. $p=1-F_{S}(s;\ell(i+1))$, selecting the $m$ lowest $p$-values amounts to keep the $m$ candidates maximizing the cumulative scores, hence the name of our scheme: WaterMax.
4.3 Low latency
Another issue is the latency: the final text cannot be issued until all candidates reach an end of sequence. This problem is fixed with the drastic choice $m=1$. It amounts to a greedy search appending to the text at iteration $i$ the chunk draft yielding the biggest incremental score $\delta s_{i,1,k}$. This choice enables the output of a chunk of text at the end of each iteration, hence reducing latency. Of note, this also corresponds to the draft with the lowest local $p$-value $1-F_{S}(\delta s_{i,1,k};\ell)$.
4.4 Optimal detection without attack
This section introduces a simple detector to reveal the advantage of our watermark embedding. The idea is that the received text is composed of chunks whose local $p$-values are distributed as beta distributions $B(1,1)$ (i.e. $\mathcal{U}_{[0,1]}$) under $\mathcal{H}_{0}$ or $B(1,n)$ under $\mathcal{H}_{1}$ (see Sect. 3.2). This is a generalization of the algorithm of the previous section, the only difference being that the likelihood ratio (1) now aggregates a vector $\mathbf{p}$ of $N$ observed local $p$-values:
Proposition 4.1.
The optimal detector $\Lambda_{\mathsf{opt}}$ has the following test’s performance when there is no attack:
$P_{FA}$ $=\gamma_{N,1}\left(\tau\right),$ (5)
where $\gamma_{x,y}$ is the lowest incomplete gamma function with shape parameter $x$ and scale parameter $y$.
Appendix B gives a sketch of the proof.
Again, notice that neither the distribution of the token variables nor the length of the chunks plays a role in the test performance (with the important caveat that a chunk must not be empty). Figure 3 illustrates the major impact of the number of chunks. Contrary to Sect. 3, we can easily reach a power close to $1$ even at $P_{FA}=10^{-6}$. For example, using $N=9$ chunks of $n=15$ drafts, assuming a medium-size text of $L=512$ tokens and chunks of equal length, one only needs to generate $Nn=135$ chunk drafts of $\ell=\lceil L/N\rceil=57$ tokens to obtain a power of $0.96$. This is equivalent to generating $15$ texts of size $512$. This is to compare to Sect. 3 where more than 3 million texts of $512$ tokens are necessary to reach this power at $P_{FA}=10^{-6}$.
Figure 3: Theoretical power of the optimal (4) and robust 7 tests at $P_{FA}=10^{-6}$ for $m=1$ as a function of the number of drafts $n$ per chunk and the number of chunks $N$.
5 Robust watermark detection
Until this point, our detector assumed that the text it receives is the generated text. In other words, the model of the previous section does not consider possible attacks. Indeed, before reaching the detector, a text might have been modified for legitimate or malicious reasons: it may be translated, curated, locally modified etc.
5.1 Robust detection
Detector (4) is neither optimal nor robust under attack because the insertion or the removal of tokens provokes a desynchronization: The detector no longer knows where each chunk starts and ends, which is necessary for computing the local $p$-value of individual chunks. More robustness comes with a global score summing up over all the tokens:
$\Lambda_{\mathsf{rob}}(\mathbf{u})=\sum_{i=1}^{N}\sum_{j=1}^{\ell}u_{(i-1)\ell+j}.$ (7)
Proposition 5.1.
The robust detector has the following test performance when there is no noise:
$P_{FA}$ $=$ $\Phi(-\tau/\sqrt{L}),$ (8)
$P_{D}$$\approx$ $\Phi\left(\frac{\Phi^{-1}(P_{FA})+\sqrt{N}e(n)}{\sqrt{v(n)}}\right).$ (9)
Appendix C gives a sketch of the proof.
The robust test (7) is less powerful than (4) when there is no attack. Figure 3 shows that the robust test (7) is slightly less powerful than the optimal test (4) when there is no attack. The following section shows that its power degrades smoothly in case of token insertion or removal.
5.2 A model of the robustness against attack
Despite the variety of modifications that can be performed on a watermarked text, all attacks have the same end effect: modifying the tokens and potentially the number of tokens. This amounts to modifying a proportion of the token variables to be distributed following $\mathcal{H}_{0}$ instead of $\mathcal{H}_{1}$ in the global score. Formally, the score $\Lambda_{\mathsf{rob}}(\mathbf{U})$ for a text of $L$ tokens becomes: (assuming $\alpha L$ is an integer)
where $1-\alpha$ is the proportion of modified tokens whose variables are denoted $\{\bar{U}_{i}\}$, and $\pi_{0}$ (resp. $\pi_{1}$) is mapping to the indices of the untouched tokens (resp. modified tokens). It is important to note that it does not matter if the attacked text’s size differs from the generated text’s size.
It is clear that $\bar{U}_{i}\overset{\mathsf{iid}}{\sim}\mathcal{N}(0;1)$. However, Appendix D shows that $\Lambda_{\mathsf{rob}}(\mathbf{U})$ is not distributed as $\alpha S_{1}+(1-\alpha)S_{0}$ where $S_{i}$ is the score under $\mathcal{H}_{i}$.
Proposition 5.2.
The robust detector has the following test performance under attack:
$P_{FA}$ $=$ $\Phi(-\tau/\sqrt{L}),$ (11)
$P_{D}$$\approx$ $\Phi\left(\frac{\Phi^{-1}(P_{FA})+\alpha\sqrt{N}e(n)}{\sqrt{1+\alpha^{2}(v(n)-1)}}\right).$ (12)
Appendix D gives a sketch of the proof.
In the end, the power of the test decreases smoothly with the strength $(1-\alpha)$ of the attack. Observe that, once again, the power of the test does not depend on the size of the text.
6 Independence
This section discusses the major assumptions made so far.
6.1 Token variable independence
All random variables $(U_{i})$ associated with tokens must be independent and identically distributed within a text. This assumption has an impact on the power of the test and especially on the probability of false alarm. It must hold for non-watermarked text. Otherwise, the score of a chunk (or of a full text) is not properly normalized into a $p$-value. We use the exact same technique as in (Fernandez et al., 2023) to enforce the token variable independence: the variable $U_{i}$ of the $i$-th token of a text depends not only on the secret key but also on the hashing of the current token and the $h-1$ previous token of the text (a window of size $h$). Furthermore, this $h$-gram is appended to a list. If the $h$-gram is already stored in that list, its variable is discarded for computing the score. By discarding identical $h$-grams, we ensure that for a given text, the $(U_{i})$ are always different. Contrary to (Fernandez et al., 2023), this mechanism is enforced during both the generation and the detection phases to ensure that both sides compute the same score.
6.2 Draft score independence
At the embedding, the scores of the $n$ drafts of a chunk should be independent. Correlations occur when parts of texts repeat across different drafts. For example, in a prompt asking ”Who are you?”, it is likely that many drafts of the first chunk of the response start by ”I am”. This assumption only plays a role in the power of the detector.
The inter-draft independence is tackled using the same methods. Note that the longer the window, the more likely a $h$-gram is new and the more diverse the scores. Yet, this diversity is obtained at the cost of robustness. Indeed, changing a single token in a text removes the watermark of $h$ tokens because their variables depend on the modification. One case guarantees draft score independence: when beam-search generates drafts instead of sampling (thus ensuring $n$ different texts) and the whole draft is hashed for seeding each variable token. Sadly, this creates a watermark that breaks even for a single modified token. This forces us to use a causal window, meaning that the score of each token depends only on itself and the previous tokens. This has an important weakness: at a given iteration, the $j$-th token ($j<h$) of every draft always refers to the same $h-j$ tokens since the previous chunk is identical for all these drafts. This means that the effective window size for this token is always smaller than $h$. In summary, we cannot guarantee our watermarking scheme perfectly adheres to this assumption while also being robust to attacks.
6.3 Experimental validation
For the independence of the token variables, the experiment runs the detector on 100k Wikipedia entries from 2018 for ten different keys. Since the text of Wikipedia is free of any watermark, one should observe $p$-values uniformly distributed, i.e. the empirical false alarm rate should match the theoretical probability of false alarm. Figure 4 demonstrates that this assumption holds for any window size $h$. On the other hand, the tamper resistance highly depends on the hashing window size, with $h=4$ being a good trade-off between performance and robustness.
Figure 4: Empirical $P_{FA}$ as a function of the theoretical $P_{FA}$ for $\Lambda_{\mathsf{opt}}$ on 10$\times$100k Wikipedia entries truncated at 256 tokens. Both quantities match whatever thhe window size $h$ for the hash, demonstrating the token variable independence.
Figure 5: Histograms of watermark scores for $N=16$ chunks and $n=5$ drafts. The black curve is the theoretical p.d.f.
The second experiment measures how much we deviate from the draft score independence. It generates $100$ texts of $128$ tokens on the “fake news” task from (Piet et al., 2023).The entropy of the generated texts may be low for a given chunk so that some parts of a text may be redundant among the drafts. Figure 5 reports how much our scores under $\mathcal{H}_{1}$ deviate from the theoretical distribution.
First, our theoretical model perfectly matches the empirical distribution when we use beam search and a non-causal window. This is expected since this setup guarantees the draft score independence if at least one token differs between the drafts. On the other hand, a causal window, whatever its size $h$, provokes a mismatch with the theory. Interestingly, the empirical distribution is close to the theoretical distribution for fewer drafts per chunk: the redundancy within the drafts makes a smaller “effective” $n$. Note that the impact is limited. Figure 3 shows that $n$ slightly weakens the test’s power compared to the number of chunks $N$. The watermark remains highly detectable as long as $N$ is large because the draft score dependence impacts only $n$,
7 Experiments
Figure 6: Detectability of WaterMax (optimal and robust), ‘distribution shift’, and ‘exponential’ without attack. (Left) histogram of $p$-values, (Right) ROC curve.
Figure 7: Detectability of WaterMax (optimal and robust), ‘distribution shift’ and ‘exponential’ with the MMW suite of attacks. (Left) histogram of $p$-values, (Right) ROC curve.
7.1 Metrics
The evaluation of a watermarking scheme is based on three quantities: 1) the quality of the watermarked text relative to the non-watermarked text, 2) the detectability of the watermark, and 3) its robustness against attacks.
Text quality Two metrics evaluate the quality of watermarked text: 1) the rating given by Llama2 when prompted to evaluate the text as proposed by Piet et al. (2023) in their Mark My Words (MMW) benchmark and 2) the measure of perplexity proposed by Kirchenbauer et al. (2023a) using OPT-2.7b (Zhang et al., 2022) on the newslike C4 dataset (Raffel et al., 2020) .
Detectability Piet et al. (2023) measure watermark detectability by the watermark size defined in MMW. This is the median number of tokens necessary to detect the watermark at $P_{FA}=0.02$. We also report the $p$-values and ROC curves for each benchmarked dataset in Figure 6-7.
Robustness The MMW benchmark is the first to propose a full suite of attacks on the watermarked texts, including swap (remove, add, or swap words), replacement with synonym, paraphrasing, translation, contraction, expansion, lower case, misspelling, and typos addition. See (Piet et al., 2023) for details. The measure, named tamper resistance, aggregates these results.
All quantities related to the MMW benchmark are measured with the code provided by the authors222github.com/wagner-group/MarkMyWords.
7.2 State-of-the-art methods
The benchmark MMW considers four techniques names as ‘binary’ (Christ et al., 2023), ‘distribution shift’ (Kirchenbauer et al., 2023a), ‘exponential’ (Aaronson & Kirchner, 2023), ‘inverse transform’ (Kuditipudi et al., 2023). It concludes that today’s best methods are ‘distribution shift’ and ‘exponential’. This is visible in Fig. 1 where these two methods are the only ones approaching the upper left corner. We set the green-list ratio of ‘distribution shift’ to $\gamma=0.5$ following the recommendation of Piet et al. (2023).
All the evaluations use Llama2-7b-chat (Touvron et al., 2023) with temperature $T=1$. The length of a chunk $\ell$ is fixed a priori to $L/N$ where $L$ is the maximum number of tokens allowed in the benchmark (1024 for MMW and 200 for C4) and $N$ the number of chunks. The LLM may generate texts shorter than $L$ tokens, leading to less than $N$ chunks, and the last chunk may not have $\ell$ tokens exactly. Except when specified otherwise, the watermark methods all use a window size of $4$ for hashing.
7.3 Results
Figure 8: Tamper resistance as function of watermark size at $P_{FA}=0.02$. The configurations leading to a relative quality of less than $0.9$ and relative perplexity less than $0.8$ are shown transparent. The values presented here correspond to the configurations of the watermarks in Figure 6.
Figures 1 summarizes the Mark My Words benchmark outcomes. They confirm that WaterMax produces detectable watermarked texts whose quality matches the one of the original LLM output, even for a small watermark size as low as $32$ tokens (with $(N,n=(64,14)$). Figure 9 draws similar conclusions. The perplexity increases slowly but tends to stay close to the original one, whatever the detectability of the watermark. This is in stark contrast to ‘distribution shift’ reaching lower $p$-values at the cost of a greatly increased perplexity. Importantly, the relative perplexity of ‘exponential’ fastly degrades since one modifies the sampling temperature to vary the watermark strength.
A close look at Fig. 6 validates the greater impact of the number of chunks $N$ compared to the number of drafts $n$. Indeed, with $N=64$ for MMW, we reach a power close to $1$ at $P_{FA}=10^{-10}$ at virtually no quality loss for $n\in(8,14)$. In comparison, ‘distribution shift’ at $\delta=0.25$ has an equivalent power for $P_{FA}=10^{-4}$, yet at a higher cost to both perplexity and Llama2 rating.
The robust detector $\Lambda_{\mathsf{rob}}$ loses almost no power compared to $\Lambda_{\mathsf{opt}}$ when no attack is performed, while being more robust against the MMW attack suite, see Figure 7. Furthermore, our scheme also reaches higher tamper-resistance than other watermark with similar quality as shown in Fig. 8. Theoretically, Eq. (12) tells that we can reach an arbitrary high tamper-resistance by increasing the number of chunks. This should compensate for a small $\alpha$, i.e. a strong attack. Yet, the computational cost of doing may become prohibitive: WaterMax strikes a trade-off between complexity and robustness when the other schemes trades either quality or watermark size for robustness. All their configurations maximizing tamper-resistance in Fig. 8 lead to a tangible loss of quality (relative quality lower than $0.9$, relative perplexity lower than $0.8$).
Figure 9: Median of the $\log_{10}p$-value as a function of the relative perplexity measured by the oracle model OPT-2.7b. The relative perplexity is computed on texts having the same size ±$\pm 10\%$. The values presented here correspond to the configurations of the watermarks in Figure 6
7.4 Computational complexity
At face value, the computational cost of our scheme is nothing more than $Nn$ text generations, the computation of the scores being negligible. However, the cost of adding one more chunk is higher than the cost of generating one more draft. Indeed, the $n$ drafts of a chunk are sampled independently (or by a beam search), making their computation highly parallelizable on modern GPUs. On the other hand, we cannot parallelize the computations over the chunk. Since $N$ has the highest impact on the power of WaterMax, it is the main limiting factor for its performance. See Appendix E for experimental running times.
8 Conclusion
Contrary to previous art, the design of WaterMax starts from a detector and then constructs the generator to maximize the detection power. Consequently, this watermark enjoys several appealing properties: it reaches both high quality and robustness even on short texts and it does not modify any component of the LLM in any way. This is at the cost of computational complexity, which stays limited thanks to the parallelization possibilities of modern GPUs. This design complies with common tricks in the literature, like adapting the sampling temperature (Idrissi et al., 2023) or embedding short messages (Fernandez et al., 2023).
9 Impact statement
This paper presents work whose goal is to improve methods for the identification of texts generated by LLMs. Societal consequences are mainly beneficial, increasing the transparency and traceability of AI as required in recent legislations (NIST, 2023; EUParliament, 2023; CAC, 2022).
Appendix A Possible detection schemes
  • The received text is decomposed into a sequence of, say $L$, tokens
  • The score is the sum of the variable: $s=\sum_{i=h+1}^{L}u_{i}$
  • The score is converted into a $p$-value: $p=1-F_{S}(s;L)$.
The watermark embedding methods are not detailed because our scheme does not make use of them. This summary is inspired from the work of Fernandez et al. (2023).
For the detection of (Kirchenbauer et al., 2023b), the output of the PRNG is used to create a ‘green’ list which is a subset of $\mathcal{V}$ of size $\gamma|\mathcal{V}|$. Variable $U_{i}$ is set to 1 if the $i$-th token belongs to this list, otherwise $U_{i}=0$. Therefore, under $\mathcal{H}_{0}$, $U_{i}\overset{\mathsf{iid}}{\sim}\mathcal{B}(\gamma)$ and $S$ follows a binomial distribution $\mathcal{B}(L-h,\gamma)$. This makes
$p=I_{\gamma}(s,L-h-s+1),$ (13)
where $I$ is the regularized incomplete Beta function.
For the detection of (Aaronson & Kirchner, 2023), the output of the PRNG is used to draw $U_{i}\overset{\mathsf{iid}}{\sim}\mathcal{E}(1)$ so that $S$ is Gamma distributed under $\mathcal{H}_{0}$. This makes:
$p=\frac{\Gamma(L-h,s)}{\Gamma(L-h)},$ (14)
where $\Gamma$ is the upper incomplete gamma function. This formulation also holds for (Christ et al., 2023) with the reduction to binary alphabet $|\mathcal{V}|=2$.
For the detection of (Kuditipudi et al., 2023), the output of the PRNG is used to draw $U_{i}$ whose distribution matches the distribution of the product of two independent r.v. $\mathcal{U}([-1/2,1/2])$, i.e. $F_{U}(u)=2u(1-\log(4|u|))+1/2$. There is no closed form to turn the global score into a $p$-value. The authors use either a bound (see Lemma 2.4) or an empirical $p$-value computed by a costly Monte Carlo simulation with random seeds. This is the reason why this studies fixes the probability of false alarm to a rather large value, i.e. $10^{-2}$.
$p=\Phi\left(-\frac{s}{\sqrt{L}}\right).$ (15)
Appendix B Proof of proposition 4.1
The received text is composed of chunks whose local $p$-values are distributed as beta distributions $B(1,1)$ (i.e. $\mathcal{U}_{[0,1]}$) under $\mathcal{H}_{0}$ or $B(1,n)$ under $\mathcal{H}_{1}$ when there is no attack (see Sect. 3.2). Knowing that if $X\sim B(1,\beta)$ then $-\log(1-X)\sim\mathcal{E}(\beta)$, $\Lambda(\mathbf{P})$ is the sum of independent exponential r.v. and thus follows a Gamma distribution under both hypothesis but with different parameters:
This leads to the characterization of the test’s performance:
$P_{FA}$ $=\gamma_{N,1}\left(\tau\right),$ (18)
Appendix C Proof of proposition 5.1
Under $\mathcal{H}_{1}$, $\Lambda_{\mathsf{rob}}(\mathbf{U})$ can be written as the sum of $N$ independent chunk scores: $\Lambda_{\mathsf{rob}}(\mathbf{U})=\sum_{i=1}^{N}M_{i}^{(n)}$. These chunk scores are distributed as the maximum of $n$ Gaussian variables $\mathcal{N}(0;\ell)$, corresponding to the score of the drafts. Those maxima have the c.d.f. $F_{M^{(n)}}(x)=\Phi\left(x/\sqrt{\ell}\right)^{n}$ and one can compute numerically the expectation $\mathbb{E}(M^{(n)})=\sqrt{\ell}e(n)$ and the variance $\mathbb{V}(M^{(n)})=\ell v(n)$, see Tab. 1.
$\mu_{1}$ $=$ $N\sqrt{\ell}e(n)=\sqrt{NL}e(n),$ (20)
$\sigma_{1}^{2}$ $=$ $N\ell v(n)=Lv(n).$ (21)
This gives the following power of the test:
$n$ 1 2 3 4 5 6 7 8 9
$e(n)$ 0 0.56 0.84 1.03 1.16 1.26 1.35 1.42 1.48
$v(n)$ 1 0.68 0.56 0.49 0.45 0.42 0.40 0.37 0.36
Appendix D Proof of proposition 5.2
$\left(\mathbf{1}_{\alpha L}^{\top}\mathbf{A}\mathbf{U}\mid\sum_{i=0}^{L}U_{i}=s\right)\sim\mathcal{N}\left(\alpha S;L\alpha(1-\alpha)\right)$ (24)
According to App. C, an approximation for $N$ large enough is $S\sim\mathcal{N}(\mu_{1};\sigma_{1}^{2})$. The law of total variance gives
Appendix E Experimental running time
Figure 10: Running time in seconds of WaterMax as a function of the number of chunks $N$ and the number of drafts/chunk $n$ for texts of $64$ tokens using sampling on a Nvidia A40.
Appendix F Further experimental results
(a) WaterMax $n=14$
(b) WaterMax $n=14$ (Zoomed in)
(c) Distribution shift
(d) Exponential
Figure 11: Relative perplexity of each of the studied watermarks using OPT-2.7b as the oracle model. Note that perplexity was constant in $n$ for WaterMax, hence we only provide results for the highest $n$ studied.
(a) WaterMax $N=64$
(b) Distribution shift
(c) Exponential
Figure 12: MAUVE score as a function of watermark parameters for two different LLMs.
For all experiments in these supplementary materials, we used the three MMW tasks as defined in (Piet et al., 2023),namely:

### v2

@@ Appendix F Pseudo-code of the watermark embedding
Appendix G Experimental running time
Figure 7: Running time in seconds of WaterMax as a function of the number of chunks $N$ and the number of drafts/chunk $n$ for texts of $256$ tokens using sampling on a Nvidia A100.
We report the running time averaged on $5$ texts of $256$ tokens generated using Llama-2-chat-hf at temperature $1.0$ and nucleus sampling $(top_{p}=0.95)$ for different parameters in Figure 7.
Overall, generating the $296$ texts of the ’Mark My Words’ benchmark on a single A100, with maximum batch-size, takes approximately $6$min for Aaronson and KGW. For WaterMax at $(n,N,b)=(10,16,6)$, it takes approximately $30$min. The resource usage report of the cluster used to run the experiments, taking into account both CPU and GPU cores, indicates a total use of 5024 core hours.
Appendix H WaterMax is almost distortion-free
This appendix elaborates a theoretical model to justify why WaterMax is close to distortion-freeness.
Consider a given iteration of WaterMax where one particular draft among $n$ is selected as the output chunk. Denote the set of possible chunks by $\mathcal{C}$. If the chunks are all composed of $\ell$ tokens, then $|\mathcal{C}|=|\mathcal{V}|^{\ell}$. Each chunk $c_{i}\in\mathcal{C}$ has a probability $p_{i}$ to be sampled by the LLM. We denote by $\tilde{p}_{i}$ the probability that WaterMax selects this chunk. This probability is computed on expectation over the set of secret keys.
A watermarking scheme is distortion-free if $\tilde{p}_{i}=p_{i},\,\forall p_{i}\in[0,1]$. This appendix shows this is not the case for WaterMax. Yet, since we operate on chunks, i.e. sequences of tokens, and not individual tokens, their probabilities are small in practice. In this regime, we show that $\tilde{p}_{i}\approx p_{i}$, which theoretically justifies the high text quality.
H.1 $n$ draws with replacement
This version of WaterMax samples $n$ drafts in $\mathcal{C}$. This corresponds to draws with replacement because a draft can be sampled several times.
Denote by $X_{i}$ the number of times the $i-th$ chunk is drawn over $n$ samples. It follows the binomial distribution $X_{i}\sim\mathcal{B}(n,p_{i}),\,\forall i\in\{1,\ldots,|\mathcal{C}|\}$.
Denote by $Y_{i}$ the binary random variable s.t. $Y_{i}=1$ if $X_{i}>0$. It follows the Bernoulli distribution $Y_{i}\sim\mathcal{B}(1-(1-p_{i})^{n}),\,\forall i\in\{1,\ldots,|\mathcal{C}|\}$.
Denote by $\mathcal{S}$ the subset of chunks which have been sampled, i.e. $\mathcal{S}=\{c_{i}:Y_{i}=1\}\subset\mathcal{C}$. The size of this set is a random variable ranging from 1 to $n$ and whose expectation is $\mathbb{E}(|\mathcal{S}|)=\sum_{i=1}^{|\mathcal{C}|}Y_{i}=|\mathcal{C}|-\sum_{i=1}^{|\mathcal{C}|}(1-p_{i})^{n}$.
For a given watermark secret key, i.e. a given instance of the token variables $\{U_{j}\}_{j=1}^{|\mathcal{V}|}$, WaterMax computes the score of each draft in $\mathcal{S}$, and it chooses the one with the maximum score. Over the ensemble of secret keys, there is no reason why a given draft is more likely to be chosen due to the symmetry of the random token variables. Therefore, a draft in a given subset $\mathcal{S}$ has a probability $1/|\mathcal{S}|$ to be chosen on expectation over the secret keys. This yields the following global probability:
$\tilde{p}_{i}=\sum_{\mathcal{S}:c_{i}\in\mathcal{S}}\frac{1}{|\mathcal{S}|}\mathbb{P}(\mathcal{S}).$ (26)
Upper bound
The above summation includes the subset $\mathcal{S}=\{c_{i}\}$ which happens with probability $p_{i}^{n}$. For all the other subsets, which include $c_{i}$, $|\mathcal{S}|\geq 2$. This gives an upper bound:
$\tilde{p}_{i}\leq UB(p_{i})=p_{i}^{n}+\frac{1}{2}\left(1-(1-p_{i})^{n}-p_{i}^{n}\right)=\frac{1}{2}\left(1-(1-p_{i})^{n}+p_{i}^{n}\right).$ (27)
Lower bound
The above summation can be written as $\tilde{p}_{i}=\sum_{k=1}^{n}\sum_{\mathcal{S}:X_{i}=k}\frac{1}{|\mathcal{S}|}\mathbb{P}(\mathcal{S})$. When $X_{i}=k$, there are at most $n-k$ other unique sampled chunks: $|\mathcal{S}|\leq n-k+1$. This makes the following lower bound:
$\tilde{p}_{i}\geq LB(p_{i})=\sum_{k=1}^{n}\frac{1}{n-k+1}\binom{n}{k}p_{i}^{k}(1-p_{i})^{n-k}=\frac{1-(1-p_{i})^{n+1}-p_{i}^{n+1}}{(n+1)(1-p_{i})}.$ (28)
Remarks
Note that $\tilde{p}_{i}=p_{i},\,\forall p_{i}\in[0,1]$ when $n\in\{1,2\}$ because the lower and upper bounds correspond to the identity. The choice $n=1$ is not an option since it corresponds to the regular use of the original LLM without watermarking. The choice $n=2$ allows watermarking with the distortion-free property. The price to pay is a low robustness.
For $n>2$, WaterMax is not distortion-free. Indeed the upper bound (27) is lower than $p_{i}$ for $p_{i}>1/2$ showing that $\tilde{p}_{i}$ does not equal $p_{i}$ on the interval $(1/2,1)$. However, note that
$\lim_{p_{i}\to 0}\frac{LB(p_{i})}{p_{i}}=1.$ (29)
This property explains why WaterMax reaches text quality as good as other distortion-free watermarking schemes. WaterMax works with chunks of text whose probabilities are all very small, contrary to other schemes working on tokens which sometimes have a picky distribution (one token gets almost all the probability mass). Therefore, missing the distortion-free property for large $p_{i}$ is irrelevant because this case never happens in practice when dealing with chunks. On the contrary, having $\tilde{p}_{i}\approx p_{i}$ locally for $p_{i}\to 0$ is key.
H.2 $n$ draws without replacement
This version of WaterMax samples $n$ drafts in $\mathcal{C}$ making sure that they are all different thanks to the beam-search suggested in Sec. 6.2. This corresponds to draws without replacement because a draft cannot be sampled twice. In other words, $|\mathcal{S}|=n$. Equation (26) simplifies to:
$\tilde{p}_{i}=\frac{1}{n}\mathbb{P}(c_{i}\in\mathcal{S}).$ (30)
However, the probability $\mathbb{P}(c_{i}\in\mathcal{S})$ is cumbersome to calculate as it relates to the multinomial Wallenius noncentral hypergeometric distribution. When all the chunk probabilities are very small, one can show that [12]:
$\tilde{p}_{i}\approx\frac{p_{i}}{n}\sum_{k=0}^{n-1}(1-p_{i})^{k}=\frac{1-(1-p_{i})^{n}}{n}\overset{p_{i}\to 0}{\longrightarrow}p_{i}.$ (31)
The same rationale as above applies: WaterMax is not distortion-free but tends to this property when all the chunks have small probabilities. This is what happens in practice when working with long chunks.
Appendix I Impact of hashing window size and beam-search tokens
Section 6 discusses how the window hashing size $h$ as well the number of tokens generated by beam-search $b$ can help to increase detectability. Both parameters, however, lead to a trade-off. First, a higher $h$ provides better detectability when no attack is performed on the text but makes the watermark less robust since modifying one token leads to at most $h$ scores, which become unusable. Secondly, a lower $b>0$ leads to better detectability by enforcing diversity between chunk drafts, at no cost to robustness, but at the cost of text quality since the beam-search has more chance of selecting tokens in the tail of the distribution. Note that $b=0$ corresponds to the baseline WaterMax, which is empirically lossless.
To select the best trade-off, we generate $296$ texts of $256$ tokens following the same protocol described in Sect. 6: we use Llama-3-8b-Instruct with temperature $\theta=1.0$, nucleus sampling ($top_{p}=0.95$), and WaterMax at $(n,N)=(8,16)$. We report the results in Fig. 9-10.
From Figure 9-9, we see that the most significant improvement in the trade-off detectability vs. quality is by going from $b=0$ to $b=6$. This is done at almost no cost to the text quality. A larger gain in detectability is obtained for $61\leq b<$ 6 but with a significant loss of quality.
Regarding the window size, $h=2$ leads to a weak scheme. An acceptable performance, close to $P_{D}=1.0$ at $P_{FA}=10^{-6}$ without attack, requires $h\geq 4$. However, Figure 10 reveals that a close match to the theoretical score distribution under $\mathcal{H}_{1}$ is only reached for $h=6$.
Figure 8: Detectability as a function of the number of tokens generated by beam-search $b$
Figure 9: Detectability as a function of the text quality. The points correspond to the different $b$ used in Fig. 9
(a) $h=2$
(b) $h=4$
(c) $h=6$
Figure 10: Score independence as a function of the window hashing size $h$ and number of tokens generated by beam-search $b$.
Appendix J Impact of the LLM and its entropy
This appendix studies the impact of the LLM on the watermark performance. As we alluded to numerous times, the performance of the watermark of KGW and Aaronson is highly dependent on the entropy of the completion of individual tokens. This dependence is well-illustrated by the impact of temperature on their performance. However, this also means that we can expect vastly different behaviors of these schemes depending on the LLM used since, depending on how the LLM was trained and fine-tuned, the average entropy of the completions varies wildly.
To demonstrate this point, we repeated the experiments in Sect. 7 for two other models: the older Llama-2-7b-chat [31] and the smaller Phi-3-mini-4k-Instruct [2] (3.8 billion parameters). We report the results, along with those of Llama-3-8b-Instruct in Figures 12-14.
The performances of both KGW and Aaronson are abysmal on Llama-2-7b-chat, with extremely low detectability, even for a high $\delta$ for KGW. At the opposite end, every algorithm performs well on Phi-3-mini-4k-Instruct. However, once again, KGW still incurs a high cost in terms of text quality, whereas WaterMax $b=6$ and Aaronson are both virtually lossless at this regime. Llama-3-8b-Instruct sits in between these two extremes.
To understand these results and verify our claim with regard to entropy, we generated $64\times 100$ completions using each model. Fig. 11 reports the distribution of their entropy. Very clearly, the entropy follows our observation, with Phi-3-mini-4k-Instruct having the highest average entropy whereas Llama-2-7b-chat has the lowest with a lot more tokens which are close to being deterministic.
This is an essential observation for watermark evaluation since different schemes are better suited to different LLMs. However, WaterMax is currently the only scheme that consistently attains almost perfect detectability at $P_{FA}=10^{-6}$ irrespective of the LLM, of the temperature, and at almost no cost to text quality.
Figure 11: Entropy of the distribution of $64\times 100$ completions for different LLM at temperature $\theta=1.0$ using nucleus sampling ($top_{p}=0.95$).
(a) Detectability as a function of the temperature of the LLM.
(b) Detectability as a function of relative perplexity. Each point corresponds to one temperature in the left figure.
Figure 12: Detectability against quality with Llama-2-7b-chat-hf, nucleus sampling ($top_{p}=0.95$), and hashing window $h=6$.
(a) Detectability as a function of the temperature of the LLM.
(b) Detectability as a function of relative perplexity. Each point corresponds to one temperature in the left figure.
Figure 13: Detectability against quality with Llama-3-8b-Instruct, nucleus sampling ($top_{p}=0.95$), and hashing window $h=6$.
(a) Detectability as a function of the temperature of the LLM.
(b) Detectability as a function of relative perplexity. Each point corresponds to one temperature in the left figure.
Figure 14: Detectability against quality with Phi-3-mini-4k-instruct, nucleus sampling ($top_{p}=0.95$), and hashing window $h=6$.
Appendix K Further experimental results
This section reports more results of interest to the practitioner to tune their watermark of choice. In particular, we report different quality metrics by varying the watermarking parameters. We also study the impact of text size on the detectability of watermarks.
In this appendix, the experiments are based on three MMW tasks defined in [28], namely: